OpenStudy (anonymous):

Find the inverse of f(x)=\[\frac{ 1 }{ 1-x }\]

OpenStudy (agent47):

y=1/(1-x) solve for x: (1-x)y=1 1-x=1/y -x=(1/y)-1 x=1-(1/y)

OpenStudy (agent47):

Now just replace x by f(x) and y by x, and voila, you're done.